If a:b and b:c = 4/3, than c:a = ? ?
Thank you.

Answer by TimothyLamb(4379)

a/b = 4/3
b/c = 4/3
---
a = (4/3)b
c = (3/4)b
---
c/a = (3/4)b/(4/3)b
c/a = (3/4)/(4/3)
c/a = (3/4)*(3/4)
c/a = 9/16
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---